Island Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each product’s selling price, variable expense per unit, and annual sales volume are as follows:

Hawaiian Fantasy Tahitian Joy
Selling price per unit $ 30 $ 125
Variable expense per unit $ 21 $ 25
Number of units sold annually 10,000 5,600

Fixed expenses total $565,500 per year.

Required:

1. Assuming the sales mix given above, do the following:

a. Prepare a contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

b. Compute the company's break-even point in dollar sales. Also, compute its margin of safety in dollars and its margin of safety percentage.

2. The company has developed a new product called Samoan Delight that sells for $50 each and that has variable expenses of $35 per unit. If the company can sell 20,000 units of Samoan Delight without incurring any additional fixed expenses:

a. Prepare a revised contribution format income statement that includes Samoan Delight. Assume that sales of the other two products does not change.

b. Compute the company’s revised break-even point in dollar sales. Also, compute its revised margin of safety in dollars and margin of safety percentage.
Answer #1

1.

a.

Hawaiian Fantasy Tahitan joy Total
10000 units 5600 units
Sales 30 300000 125 700000 1000000
Less: Variable cost 21 210000 25 140000 350000
Contribution 9 90000 100 560000 650000
% contribution of sales 30% 80% 65%
Less : Fixed costs 565500
Margin 90000 560000 84500

b.

Break Even Sales: Fixed cost/Contribution margin => 565500/0.65 => 870000.

Margin of Safety: 1000000-870000 = 130,000. So, 130000/1000000*100 = 13%.

2.

a.

Hawaiian Fantasy Tahitan joy Samoan Delight Total
10000 units 5600 units 20000 units
Sales 30 300000 125 700000 50 1000000 2000000
Less: Variable cost 21 210000 25 140000 35 700000 1050000
Contribution 9 90000 100 560000 15 300000 950000
% contribution of sales 30% 80% 30% 47.5%
Less : Fixed costs 565500
Margin 90000 560000 300000 384500

b.

Break Even Sales: Fixed cost/Contribution margin => 565500/0.475 => 1190526.

Margin of Safety: 2000000-1190526 = 809474. So, 40.44%